Carter Company's return on common equity is 29%. Its sales are $68,000,000, its debt ratio is 45%, and its total liabilities are $15,000,000. What is the firm's return on total assets?

Question 4 options:

15.95%

13.05%

17.61%

21.20%

10.87%
Answer #1
Debt ratio = 45%

It means total capital or Total assets = 1

So, Equity = 1-45% = 55%

Total libilities = 15000000

So, Equity = Total liabilities / Debt ratio * Equity

15000000/45% * 55%

18333333.33

Total assets = Total liabilities / Debt ratio *1

33333333.33

Return on equity = 29%

So, Net income or return = Equity * ROE

18333333.33*29%= 5316666.667

Return on Total assets = Net income / Total assets

5316666.667 / 33333333.33
0.1595 or 15.95%

So, return on total assets is 15.95%

Please thumbs up.